OpenStudy (anonymous):

your sock drawer holds 2 blue, 5 tan, and 4 brown socks. without looking, you grad one sock. find the probability that the sock is brown.

OpenStudy (anonymous):

4/11

OpenStudy (anonymous):

is it 4/11

OpenStudy (callisto):

P(brown) = no. of brown / total = 4 / (2+5+4) = 4/11